# Global Bifurcation of Non-Radial Solutions for Symmetric Sub-linear Elliptic Systems on the Planar Unit Disc

### v1

@@ Abstract.
In this paper, we prove a global bifurcation result for the existence of non-radial branches of solutions to the paramterized family of $\Gamma$-symmetric problems $-\Delta u=f(\alpha,z,u)$, $u|_{\partial D}=0$ on the unit disc $D:=\{z\in{\mathbb{C}}:|z|<1\}$ with $u(z)\in\mathbb{R}^{k}$, where $\mathbb{R}^{k}$ is an orthogonal $\Gamma$-representation, $f:\mathbb{R}\times\overline{D}\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ is a sub-linear $\Gamma$-equivariant continuous function, differentiable with respect to $u$ at zero and satisfying the conditions $f(\alpha,e^{i\theta}z,u)=f(\alpha,z,u)$ for all $\theta\in\mathbb{R}$ and $f(z,-u)=-f(z,u)$.
@@ 1. Introduction
We draw our motivation from natural phenomena that are modeled by parameterized systems of autonomous Partial Differential Equations (PDEs), where the possible states of each phenomenon are represented by solutions of the associated system of equations at corresponding parameter values. These solutions can be categorized as trivial or non-trivial. Trivial solutions are so-named because they persist for all parameter values and are straightforward to determine. In contrast, Non-trivial solutions may only exist for a particular range of parameter values and often exhibit exotic properties which enhance our understanding of the model. The classical bifurcation problem is concerned with the existence and global behaviour of branches of non-trivial solutions emerging from their trivial counterparts. Should the phenomena admit the symmetries of a certain group $G$ (represented by the $G$-equivariance of the associated equations), one may study the equivariant bifurcation problem, which additionally considers the symmetric properties of these branches (cf. [1, 3, 9, 11, 19, 21, 22, 23, 24, 25]). The application of topological methods to the study of differential equations, as with most tools in the arsenal of the nonlinear analyst, traces back to Maître Poincare [13]. More recently, the Local/Equivariant Brouwer Degrees and their infinite dimensional generalizations – the Local/Equivariant Leray-Schauder Degrees – have proven prodigiously successful in obtaining existence results for solutions to a wide class of nonlinear differential equations. Somewhat remarkably, these degree theories have also been shown to be useful in solving classical and equivariant bifurcation problems. The first use of the Leray-Schauder degree for the detection of (local) bifurcation to a parameterized system of nonlinear differential equations is attributed to the seminal work of M.A. Krasnosel’skii, in which sufficient conditions for the existence of a branch of non-trivial solutions are established (cf. [17]). Only a decade later, P. Rabinowitz (cf. [20]) proposed his famous Rabinowitz alternative, which provides sufficient conditions for a given branch of non-trivial solutions to be unbounded. In this paper, equivariant analogues of Krasnosel’skii and Rabinowitz type results are employed to solve an equivariant bifurcation problem. Specifically, our objective is to study the equivariant bifurcation problem for a symmetric system of parameterized nonlinear Laplace equations subject to Dirichlet boundary conditions. The bifurcation and global behaviour of solutions for such systems has been extensively studied, due to their significance in applied mathematics and physics. In the case that the nonlinearity arises as the gradient of some potential function, the bifurcation problem can be studied as a parameterized variational problem and appropriate topological methods such as Morse theory, index theory, and the equivariant gradient degree (cf. [14]) can be administered (cf. [6, 7, 8, 10, 12, 11, 21, 22, 23, 24, 25, 26]). However, these approaches are of no use for systems which do not admit variational structure. The degree theoretic methods we employ in this paper impose no variational requirements on the equations and, as such, are broadly applicable to a wider class of problems. Consider the following parameterized family of symmetric Laplace equations:
@@
    radial solutions, which depend only on $|z|$ for any $z\in D$ (although less obvious than the trivial solution, radial solutions can often be identified using classical methods, eg. by reducing (1) to a second order ODE);
@@ 2. Functional Space Reformulation and a priori Bounds
where $s=(s_{1},s_{2})$, $|s|:=s_{1}+s_{2}\leq 2$, and $D^{s}\varphi=\frac{\partial^{|s|}\varphi}{\partial^{s_{1}}x\partial^{s_{2}}y}$. It is well known that the Laplacian operator $\mathscr{L}:\mathscr{H}\to L^{2}(D,V)$ given by
@@
It suffices to demonstrate that for any $v\in L^{q}(D,V)$ and $\alpha\in\mathbb{R}$ one has $N_{\alpha}(v)\in L^{2}(D,V)$. Indeed, combining ($A_{5}$) with the Hölder inequality, one has:
@@
where the result follows from the assumptions that $b(\alpha),a(\alpha),\|v\|_{q}<\infty$. $\square$ Notice that the operator equation
@@
is equivalent to system (1) in the sense that $u\in\mathscr{H}$ is a solution to (1) for some $\alpha\in\mathbb{R}$ if and only if $(\alpha,u)\in\mathbb{R}\times\mathscr{H}$ satisfies (6). In turn, since $\mathscr{L}$ is an isomorphism, the map $\mathscr{F}:\mathbb{R}\times\mathscr{H}\to\mathscr{H}$ given by
@@
is well-defined so that system (1) is also equivalent to the equation
@@ 3.1. The Local Bifurcation Invariant and Krasnosel’skii’s Theorem
Let $H\leq\mathcal{G}$ be a given subgroup and denote by $\mathcal{S}^{H}$ the corresponding $H$-fixed point space of non-trivial solutions. A branch of solutions $\mathcal{C}$ is said to have symmetries at least $(H)$ if $\mathcal{C}\cap\mathcal{S}^{H}\not=\emptyset$.
Take an isolated critical point $(\alpha_{0},0)\in\Lambda$ and choose $\alpha^{\pm}_{0}\in(\alpha_{0}-\varepsilon,\alpha_{0}+\varepsilon)$ with $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$, where $\varepsilon>0$ was chosen such that for all $0<|\alpha-\alpha_{0}|<\varepsilon$ the solution $(\alpha,0)\in M$ is a regular point. For convenience, we adopt the notations ±$\mathcal{F}_{\pm}(u):=\mathcal{F}(\alpha^{\pm}_{0},u)$ and $B_{\delta}:=\{u\in\mathscr{H}:\|u\|<\delta\}$. By construction, the linear operators $\mathcal{A}(\alpha^{\pm}):\mathscr{H}\rightarrow\mathscr{H}$ are non-singular; so, from assumption ($B_{2}$), there exists $\delta>0$ sufficiently small such that ±$\mathcal{F}_{\pm}^{-1}(0)\cap\partial B_{\delta}=\emptyset$ and $\mathcal{F}_{\pm}$ are $B_{\delta}$-admissibly $\mathcal{G}$-homotopic to $\mathcal{A}(\alpha^{\pm})$. Moreover, since $\mathcal{G}$ acts isometrically on $\mathscr{H}$, the ball $B_{\delta}$ is clearly $\mathcal{G}$-invariant. It follows that (±$(\mathcal{F}_{\pm},B_{\delta})$ are admissible $\mathcal{G}$-pairs in $\mathscr{H}$ and also that $\mathcal{G}\text{-deg}(\mathcal{F}_{\pm},B_{\delta})=\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{\pm}_{0}),B(\mathscr{H}))$, where $B(\mathscr{H})$ is the open unit ball in $\mathscr{H}$. We define the the local $G$-equivariant bifurcation invariant $\omega_{\mathcal{G}}(\alpha_{0})\in A(\mathcal{G})$ at the isolated critical solution $(\alpha_{0},0)\in\Lambda$ as follows,
$\omega_{\mathcal{G}}(\alpha_{0})=\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{-}_{0}),B(\mathscr{H}))-\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{+}_{0}),B(\mathscr{H})).$
The invariant $\omega_{\mathcal{G}}(\alpha_{0})$ does not depend on the choice of $\alpha^{\pm}_{0}\in\mathbb{R}$ or the radius $\delta>0$. We refer to [4] for the proof of the following local bifurcation result, which is a consequence of the equivariant version of a classical result of K. Kuratowski (cf. [18], Thm. 3, p. 170).
@@ 3.2. Global Bifurcation and the Rabinowitz Alternative
(The Rabinowitz Alternative) Suppose that $\mathcal{F}:\mathbb{R}\times\mathscr{H}\rightarrow\mathscr{H}$ is a completely continuous $\mathcal{G}$-equivariant field satisfying conditions ($B_{1}$),($B_{2}$) and ($B_{3}$). Let $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}$ be an open bounded $\mathcal{G}$-invariant set with $\partial\mathcal{U}\cap\Lambda=\emptyset$ and $(\alpha_{0},0)\in\mathcal{U}\cap\Lambda$. If $\mathcal{C}$ is a branch of nontrivial solutions to (11) bifurcating from the critical point $(\alpha_{0},0)$, then one has the following alternative:
@@ 2. $(b)$
Suppose that Theorem 3.1 is used to demonstrate the existence of a branch $\mathcal{C}$ of nontrivial solutions to (11) bifurcating from a critical point $(\alpha_{0},0)$. If, in addition, certain conditions are met such that for any open bounded $\mathcal{G}$-invariant neighborhood $\mathcal{U}\ni(\alpha_{0},0)$ with $\partial\mathcal{U}\cap\Lambda=\emptyset$ the alternative $(b)$ is impossible, then, according to Theorem 3.2, the branch $\mathcal{C}$ must be unbounded.
@@ 4. Local and Global Bifurcation of Non-radial Solutions in (1)
Returning to the functional reformulation of our original problem described in Section 2, put $G:=O(2)\times\Gamma\times\mathbb{Z}_{2}$ and notice that the Sobolev space $\mathscr{H}$ introduced in Section 2 is a natural Banach $G$-representation with respect to the isometric $G$-action $G\times\mathscr{H}\rightarrow\mathscr{H}$ given by
@@
Choose a value $\alpha\in\mathbb{R}$ such that $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism and denote by $B(\mathscr{H}):=\{u\in\mathscr{H}:\|u\|_{\mathscr{H}}<1\}$ the open unit ball in $\mathscr{H}$. Before approaching the question of bifurcation in the equation (1), we must derive a workable formula for the computation of the degree $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$. Assuming that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=1}^{r}$ is made available we denote by $\{\mathcal{V}_{j}^{-}\}_{j=1}^{r}$ the corresponding list of irreducible $\Gamma\times\mathbb{Z}_{2}$-representations, where the superscript is meant to indicate that each irreducible $\Gamma$-representation has been equipped with the antipodal $\mathbb{Z}_{2}$-action in the standard way (cf. [4]). Now, as a $\Gamma\times\mathbb{Z}_{2}$-representation, $V=\mathbb{R}^{k}$ has the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
$V=V_{1}\oplus V_{1}\oplus\cdots\oplus V_{r},$
@@
To simplify computations, we introduce an additional condition on the linearization of (1):
@@
In pursuit of a $G$-isotypic decomposition of $\mathscr{H}$, let us consider the spectrum of the Laplacian operator (2), understood in this context as an unbounded operator in $L^{2}(D;V)$. Namely, one has
@@
Then, $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
the $G$-isotypic of $\mathscr{H}$ is given by
@@
To be clear, each $G$-isotypic component $\mathscr{H}_{m,j}$ is modeled on the irreducible $G$-representation $\mathcal{V}_{m,j}:=\mathscr{H}_{m}\otimes\mathcal{V}_{j}^{-}$, where $\mathscr{H}_{m}$ denotes the $m$-th irreducible $O(2)$-representation and $\mathcal{V}_{j}^{-}$ is the $j$-th irreducible $\Gamma$-representation $\mathcal{V}_{j}$ with antipodal $\mathbb{Z}_{2}$-action.
@@
$\square$ Assume that, for a given $\alpha\in\mathbb{R}$, the operator $\mathscr{A}(\alpha):\mathscr{H}\to\mathscr{H}$ is an isomorphism. The product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix C) permits us to express $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$ in terms of a Burnside ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{n,m}^{j}(\alpha):\mathscr{E}^{j}_{nm}\rightarrow\mathscr{E}^{j}_{nm}$ of the $G$-equivariant linear isomorphism $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-subrepresentations $\mathscr{E}^{j}_{nm}$ on their respective open unit balls $B(\mathscr{E}^{j}_{nm}):=\{u\in\mathscr{E}^{j}_{nm}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@
Notice that, under assumption ($A_{0}$), one has $G\text{\rm-deg}(\mathscr{A}_{n,m}^{j}(\alpha),B(\mathscr{E}_{nm}^{j}))=(G)$ for almost all indices $m$, $n$ and $j$, so that the product (15) is well-defined. Indeed, each Leray-Schauder $G$-equivariant degree $G\text{\rm-deg}(\mathscr{A}_{n,m}^{j}(\alpha),B(\mathscr{E}_{nm}^{j}))$ is fully specified by the $\mathcal{V}_{j}^{-}$-isotypic multiplicities $\{m_{j}\}_{j=1}^{r}$ together with the real spectra of $\mathscr{A}(\alpha)$ according to formula,
@@
where $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is the basic degree (cf. Appendix C) associated with the irreducible $G$-representation $\mathcal{V}_{m,j}$ and $(G)\in A(G)$ is the unit element of the Burnside Ring. In addition, since each basic degree $\deg_{\mathcal{V}_{m,j}}$ is involutive in the Burnside ring (cf. Appendix C), one has
@@
Putting together (15), (16) and (17), we introduce some notations to keep track of the indices
@@
As is well known (see, for example, [27], p. 486), $s_{1m}>m(m+2)$, from which it follows that the set (18) is finite. Combining (18) with formulas (15) and (16), one obtains

### v2

@@ Abstract.
In this paper, we prove a global bifurcation result for the existence of non-radial branches of solutions to the paramterized family of $\Gamma$-symmetric equations $-\Delta u=f(\alpha,z,u)$, $u|_{\partial D}=0$ on the unit disc $D:=\{z\in{\mathbb{C}}:|z|<1\}$ with $u(z)\in\mathbb{R}^{k}$, where $\mathbb{R}^{k}$ is an orthogonal $\Gamma$-representation, $f:\mathbb{R}\times\overline{D}\times\mathbb{R}^{k}\to\mathbb{R}^{k}$ is a sub-linear $\Gamma$-equivariant continuous function, differentiable with respect to $u$ at zero and satisfying the conditions $f(\alpha,e^{i\theta}z,u)=f(\alpha,z,u)$ for all $\theta\in\mathbb{R}$ and $f(z,-u)=-f(z,u)$.
@@ 1. Introduction
We draw our motivation from natural phenomena that are modeled by parameterized systems of autonomous Partial Differential Equations (PDEs), where the possible states of each phenomenon are represented by solutions of the associated system of equations at corresponding parameter values. These solutions can be categorized as trivial or non-trivial. Trivial solutions are so-named because they persist for all parameter values and their existence is straightforward to determine. In contrast, non-trivial solutions may only exist for a particular range of parameter values and often exhibit exotic properties that enhance our understanding of the model. The classical bifurcation problem is concerned with the existence and global behaviour of branches of non-trivial solutions emerging from their trivial counterparts. Phenomena which admit the symmetries of a certain group $G$ (represented by the $G$-equivariance of the associated equations) may be studied as equivariant bifurcation problems, with additional consideration placed on the symmetric properties of these branches (cf. [1, 3, 9, 11, 19, 21, 22, 23, 24, 25]). The application of topological methods to the study of differential equations, as with most tools in the arsenal of the nonlinear analyst, traces back to Maître Poincare [13]. More recently, the Local/Equivariant Brouwer Degrees and their infinite dimensional generalizations – the Local/Equivariant Leray-Schauder Degrees – have proven prodigiously effective in obtaining existence results for solutions to a wide class of nonlinear differential equations. Somewhat remarkably, these degree theories are also instrumental in solving classical and equivariant bifurcation problems. The first use of the Leray-Schauder degree for the detection of (local) bifurcation in a parameterized system of nonlinear differential equations is attributed to the seminal work of M.A. Krasnosel’skii, in which sufficient conditions for the existence of a branch of non-trivial solutions are established (cf. [17]). Only a decade later, P. Rabinowitz (cf. [20]) proposed his famous Rabinowitz alternative, which provides sufficient conditions for the unboundedness of a branch of non-trivial solutions. In this paper, we employ equivariant analogues of Krasnosel’skii and Rabinowitz type results to solve an equivariant bifurcation problem. Specifically, our objective is to study the equivariant bifurcation problem for a symmetric system of parameterized nonlinear Laplace equations subject to Dirichlet boundary conditions. The bifurcation and global behaviour of solutions for such systems has been extensively studied, due to their significance in applied mathematics and physics. In the case that the nonlinearity arises as the gradient of some potential function, the bifurcation problem can be studied as a parameterized variational problem and appropriate topological methods such as Morse theory, index theory, and the equivariant gradient degree (cf. [14]) can be administered (cf. [6, 7, 8, 10, 12, 11, 21, 22, 23, 24, 25, 26]). However, these approaches are not applicable for systems which lack variational structure. The degree theoretic methods used in this paper impose no variational requirements on the equations and, as such, are broadly applicable to a wider class of problems. Consider the following parameterized family of symmetric Laplace equations:
@@
    radial solutions, which depend only on the magnitude $|z|$ of any disc element $z\in D$ (although less obvious than the trivial solution, radial solutions can often be identified using classical methods, eg. by reducing (1) to a second order ODE);
@@ 2. Functional Space Reformulation and a priori Bounds
where $s:=(s_{1},s_{2})$, $|s|:=s_{1}+s_{2}\leq 2$, and $D^{s}\varphi:=\frac{\partial^{|s|}\varphi}{\partial^{s_{1}}x\partial^{s_{2}}y}$. It is well known that the Laplacian operator $\mathscr{L}:\mathscr{H}\to L^{2}(D,V)$ given by
@@
It suffices to demonstrate that for any $v\in L^{q}(D,V)$ and $\alpha\in\mathbb{R}$ one has $N_{\alpha}(v)\in L^{2}(D,V)$. Combining ($A_{5}$) with the Hölder inequality, one has:
@@
where the result follows from $b(\alpha),a(\alpha),\|v\|_{q}<\infty$. $\square$ Notice that the equation
@@
is equivalent to system (1) in the sense that $u\in\mathscr{H}$ is a solution to (1) for some $\alpha\in\mathbb{R}$ if and only if $(\alpha,u)\in\mathbb{R}\times\mathscr{H}$ satisfies (6). In turn, the invertibility of $\mathscr{L}$ together with Lemma (2.1) imply that the map $\mathscr{F}:\mathbb{R}\times\mathscr{H}\to\mathscr{H}$ given by
@@
is well-defined. Hence, system (1) is also equivalent to the equation
@@ 3.1. The Local Bifurcation Invariant and Krasnosel’skii’s Theorem
Given a subgroup $H\leq\mathcal{G}$, denote by $\mathcal{S}^{H}$ the corresponding $H$-fixed point space of non-trivial solutions. A branch of solutions $\mathcal{C}$ is said to have symmetries at least $(H)$ if $\mathcal{C}\cap\mathcal{S}^{H}\not=\emptyset$.
Let $(\alpha_{0},0)\in\Lambda$ be an isolated critical point with a deleted $\epsilon$-neighborhood
$\{\alpha\in\mathbb{R}:0<|\alpha-\alpha_{0}|<\epsilon\},$
on which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism and choose $\alpha^{\pm}_{0}\in(\alpha_{0}-\varepsilon,\alpha_{0}+\varepsilon)$ with $\alpha^{-}_{0}\leq\alpha_{0}\leq\alpha^{+}_{0}$. Since $\mathcal{A}(\alpha^{\pm}):\mathscr{H}\rightarrow\mathscr{H}$ are non-singular, there exists a number $\delta>0$ sufficiently small such that, adopting the notations ±$\mathcal{F}_{\pm}(u):=\mathcal{F}(\alpha^{\pm}_{0},u)$, $B_{\delta}:=\{u\in\mathscr{H}:\|u\|<\delta\}$, one has
±$\mathcal{F}_{\pm}^{-1}(0)\cap\partial B_{\delta}=\emptyset,$
and $\mathcal{F}_{\pm}$ are $B_{\delta}$-admissibly $\mathcal{G}$-homotopic to $\mathcal{A}(\alpha^{\pm})$. Moreover, since $\mathcal{G}$ acts isometrically on $\mathscr{H}$, the ball $B_{\delta}$ is clearly $\mathcal{G}$-invariant. It follows, from the homotopy property of the $\mathcal{G}$-equivariant Leray-Schauder degree (cf. Appendix C), that (±$(\mathcal{F}_{\pm},B_{\delta})$ are admissible $\mathcal{G}$-pairs in $\mathscr{H}$ and also that $\mathcal{G}\text{-deg}(\mathcal{F}_{\pm},B_{\delta})=\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{\pm}_{0}),B(\mathscr{H}))$, where $B(\mathscr{H})$ is the open unit ball in $\mathscr{H}$. We call the Burnside Ring element
(13) $\omega_{\mathcal{G}}(\alpha_{0})=\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{-}_{0}),B(\mathscr{H}))-\mathcal{G}\text{-deg}(\mathcal{A}(\alpha^{+}_{0}),B(\mathscr{H})),$
the local bifurcation invariant at $(\lambda_{0},0)$. The reader is referred to [4] for proof that the invariant (13) does not depend on the choice of $\alpha^{\pm}_{0}\in\mathbb{R}$ or radius $\delta>0$, and also for the proof of the following local bifurcation result, which is a consequence of the equivariant version of a classical result of K. Kuratowski (cf. [18], Thm. 3, p. 170).
@@ 3.2. Global Bifurcation and the Rabinowitz Alternative
(The Rabinowitz Alternative) Suppose that $\mathcal{F}:\mathbb{R}\times\mathscr{H}\rightarrow\mathscr{H}$ is a completely continuous $\mathcal{G}$-equivariant field satisfying conditions ($B_{1}$)($B_{3}$) and let $\mathcal{U}\subset\mathbb{R}\times\mathscr{H}$ be an open bounded $\mathcal{G}$-invariant set with $\partial\mathcal{U}\cap\Lambda=\emptyset$. If $\mathcal{C}$ is a branch of nontrivial solutions to (11) bifurcating from the critical point $(\alpha_{0},0)\in\mathcal{U}\cap\Lambda$, then one has the following alternative:
@@ 2. $(b)$
Suppose that Theorem 3.1 is used to demonstrate the existence of a branch $\mathcal{C}$ of nontrivial solutions to (11) bifurcating from a critical point $(\alpha_{0},0)$ and that certain conditions are met such that, for any open bounded $\mathcal{G}$-invariant neighborhood $\mathcal{U}\ni(\alpha_{0},0)$ with $\partial\mathcal{U}\cap\Lambda=\emptyset$, the alternative $(b)$ is impossible. Then, according to Theorem 3.2, the branch $\mathcal{C}$ must be unbounded.
@@ 4. Local and Global Bifurcation of Non-radial Solutions in (1)
Returning to the functional reformulation of our original problem described in Section 2, consider the product group $G:=O(2)\times\Gamma\times\mathbb{Z}_{2}$ and notice that the Sobolev space $\mathscr{H}$ is a natural Banach $G$-representation with respect to the isometric $G$-action $G\times\mathscr{H}\rightarrow\mathscr{H}$ given by
@@
Before approaching the question of bifurcation in the equation (1), we must derive a workable formula for the computation of the degree $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$, where $B(\mathscr{H}):=\{u\in\mathscr{H}:\|u\|_{\mathscr{H}}<1\}$ is the open unit ball in $\mathscr{H}$ and $\alpha\in\mathbb{R}$ is any parameter value for which $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ is an isomorphism. Assuming that a complete list of the irreducible $\Gamma$-representations $\{\mathcal{V}_{j}\}_{j=1}^{r}$ is made available we denote by $\{\mathcal{V}_{j}^{-}\}_{j=1}^{r}$ the corresponding list of irreducible $\Gamma\times\mathbb{Z}_{2}$-representations, where the superscript is meant to indicate that each irreducible $\Gamma$-representation has been equipped with the antipodal $\mathbb{Z}_{2}$-action in the standard way (cf. [4]). As a $\Gamma$-representation, $V=\mathbb{R}^{k}$ is also a natural $\Gamma\times\mathbb{Z}_{2}$-representation with the $\Gamma\times\mathbb{Z}_{2}$-isotypic decomposition
$V=V_{1}\oplus V_{2}\oplus\cdots\oplus V_{r},$
@@
To simplify our computations, we introduce an additional condition on the linearization of (1):
@@
On the other hand, for every $m\in{\mathbb{N}}$ we denote by $\mathcal{W}_{m}\simeq{\mathbb{C}}$ the irreducible $O(2)$-representation equipped with the $O(2)$-action
$(\theta,z):=e^{im\theta}\cdot z,\;(\kappa,z):=\overline{z},\;z\in\mathcal{W}_{m},$
and by $\mathcal{W}_{0}\simeq\mathbb{R}$ the irreducible $O(2)$-representation with the trivial $O(2)$-action. In pursuit of a $G$-isotypic decomposition of $\mathscr{H}$, let us consider the spectrum of the Laplacian operator (2), understood in this context as an unbounded operator in $L^{2}(D;V)$. Namely, one has
@@
Clearly, one has
$\mathscr{E}_{nm}\simeq\mathcal{W}_{m}\otimes V,$
for each $(n,m)\in{\mathbb{N}}\times{\mathbb{N}}\cup\{0\}$ such that $\mathscr{H}$ admits the $O(2)$-isotypic decomposition
@@
one has
$\mathscr{E}_{nm}^{j}\simeq\mathcal{W}_{m}\otimes\mathcal{V}_{j},\;(n,m)\in{\mathbb{N}}\times{\mathbb{N}}\cup\{0\},\;j\in\{1,2,\ldots,r\}.$
Hence, $\mathscr{H}$ admits the $G$-isotypic
@@
To be clear, each $G$-isotypic component $\mathscr{H}_{m,j}$ is modeled on the irreducible $G$-representation $\mathcal{V}_{m,j}:=\mathcal{W}_{m}\otimes\mathcal{V}_{j}^{-}$.
@@
$\square$ For any regular point of equation (8) $\alpha\in\mathbb{R}$, the product property of the Leray-Schauder $G$-equivariant degree (cf. Appendix C) permits us to express $G\text{\rm-deg}(\mathscr{A}(\alpha),B(\mathscr{H}))$ in terms of a Burnside ring product of the Leray-Schauder $G$-equivariant degrees of the various restrictions $\mathscr{A}_{n,m}^{j}(\alpha):\mathscr{E}^{j}_{nm}\rightarrow\mathscr{E}^{j}_{nm}$ of the $G$-equivariant linear isomorphism $\mathscr{A}(\alpha):\mathscr{H}\rightarrow\mathscr{H}$ to the $G$-subrepresentations $\mathscr{E}^{j}_{nm}$ on their respective open unit balls $B(\mathscr{E}^{j}_{nm}):=\{u\in\mathscr{E}^{j}_{nm}\;:\;\|u\|_{\mathscr{H}}<1\}$ as follows
@@
Notice that one has $G\text{\rm-deg}(\mathscr{A}_{n,m}^{j}(\alpha),B(\mathscr{E}_{nm}^{j}))=(G)$ for almost all indices $m$, $n$ and $j$, so that the product (16) is well-defined. Indeed, each Leray-Schauder $G$-equivariant degree $G\text{\rm-deg}(\mathscr{A}_{n,m}^{j}(\alpha),B(\mathscr{E}_{nm}^{j}))$ is fully specified by the $\mathcal{V}_{j}^{-}$-isotypic multiplicities $\{m_{j}\}_{j=1}^{r}$ together with the real spectra of $\mathscr{A}(\alpha)$ according to formula,
@@
where $\deg_{\mathcal{V}_{m,j}}\in A(G)$ is the basic degree (cf. Appendix C) associated with the irreducible $G$-representation $\mathcal{V}_{m,j}$ and $(G)\in A(G)$ is the unit element of the Burnside Ring. In addition, since each basic degree is involutive in the Burnside ring (cf. Appendix C), one has
@@
Putting together (17) and (18), we introduce some notations to keep track of the indices
@@
As is well known (see, for example, [27], p. 486), one has
$s_{1m}>m(m+2),\;m\geq 0,$
from which it follows that the set (19) is finite. Combining (19) with formulas (16) and (17), one obtains